# Elevating All Zero-Shot Sketch-Based Image Retrieval Through Multimodal Prompt Learning

### v1

@@ 3.5 Loss functions and inference
$\mathcal{L}_{triplet}$: Cross-visual modality triplet loss with proposed adaptive margin: Given the triplet of embeddings $(\mathcal{F}_{v}(s_{a}),\mathcal{F}_{v}(p_{a}^{+}),\mathcal{F}_{v}(p_{a}^{-}))$, we formulate a triplet objective aimed at minimizing the distance between $\mathcal{F}_{v}(s_{a})$ and $\mathcal{F}_{v}(p_{a}^{+})$ while maximizing the distance between $\mathcal{F}_{v}(s_{a})$ and $\mathcal{F}_{v}(p_{a}^{-})$. Traditional triplet objectives employ a fixed margin, which may not be optimal for zero-shot tasks. In contrast, we propose leveraging the semantic space of CLIP to define the margin, utilizing the embeddings of positive and negative class names from $\mathcal{F}_{t}$. Let $\mathcal{F}_{t}(c^{+})$ and $\mathcal{F}_{t}(c^{-})$ represent the semantic class embeddings for $p_{a}^{+}/s_{a}$ and $p_{a}^{-}$, respectively. We define $\mu(c^{+},c^{-})=1-\cos(\mathcal{F}_{t}(c^{+}),\mathcal{F}_{t}(c^{-}))$. Consequently, the margin increases when $c^{+}$ and $c^{-}$ are semantically close. The loss is,
@@
$\mathcal{L}_{class}$: Text-image classification loss: A supervised contrastive loss is introduced to correctly classify sketches and photos in $\mathcal{G}^{s}$ based on the class-wise text prompts outlined in Section 3.1. In Eq. 7, $(\mathbf{I,\mathbf{Y}})$ denotes a sketch or a photo along with the associated one-hot label vector: $\mathbf{I}\in\{\mathcal{P}^{s},\mathcal{S}^{s}\}$, and $\mathcal{Y}=[y^{1},\cdots,y^{|\mathcal{C}^{s}|}]$.
@@ 4 Experiments and Results
- Datasets: We evaluate SpLIP on three benchmark datasets: Sketchy-extended [27], TU-Berlin-extended [13], and QuickDraw-extended [7], following the established training-validation protocols [5, 11]. For cross-category FG-ZS-SBIR, which necessitates precise sketch-photo matching, we incorporate the Sketchy dataset [44]. Further details are mentioned in the Supplementary.
- Implementation details, training and evaluation protocols: For $\mathcal{F}_{v}$, we select the ViT-B/32 backbone of CLIP, while the Transformer-based text encoder is considered for $\mathcal{F}_{t}$. Besides, $\mathcal{B}_{t}$ employs three linear layers to convert $\mathbf{E}_{0}$ into $m=4$ learnable textual tokens. In contrast, $\mathcal{B}_{v}$ utilizes a singular linear layer to adapt to the visual dimension, producing four learnable visual tokens for a matching batch size. Meanwhile, $\mathcal{B}_{vt}$ incorporates linear layers and employs a bottleneck architecture consisting of two layers (Linear-ReLU-Linear) for the creation of $n=2$ layer-specific visual tokens. Additionally, $\mathcal{F}_{js}$ is designed with two encoder layers, followed by a classifier, to accurately decode the patch arrangements of $s^{\prime}_{a}$ given ($p_{a}^{+}$/$p_{a}^{-}$/$s_{a}^{+}$).
The training process spans 60 epochs, initiating with a warm-up learning rate of $0.001$ and utilizing the Adam optimizer [22] alongside a scheduler. The batch size is configured to $192$ for both Sketchy and TU Berlin datasets, whereas for QuickDraw, a batch size of $64$ is adopted. $\alpha$ and $\beta$ are fixed through grid search.
@@
Following the literature [7, 55], our evaluation for ZS-SBIR considers the top $200$ retrieved photos, where we report the mean Average Precision score (mAP@all) and precision at $200$ (P@200). Aligning with the recent trends, however, we specifically report precision at $100$ (P@100) for the TU-Berlin dataset and mAP at $200$ (mAP@200) for the Sketchy dataset. For FG-ZS-SBIR, accuracy is evaluated by considering only a single category at a time [33], denoted as Acc@$\mathcal{K}$ for Sketchy. This metric reflects the percentage of sketches that have their true matched photo within the top-$\mathcal{K}$ list, with our focus being on the Top-1 and Top-5 accuracy metrics [39].
@@ 4.1 Comparison to the literature
- (G)ZS-SBIR and FG-ZS-SBIR: Table 1 provides a comparative analysis of SpLIP against methods utilizing ConvNet, ViT, and CLIP backbones across two splits of the Sketchy, TU-Berlin, and QuickDraw datasets, respectively. Remarkably, SpLIP outperforms all models based on ConvNet and ViT backbones by a substantial margin across all metrics. For instance, on the challenging QuickDraw dataset, SpLIP surpasses the leading ConvNet-based method, BDA [5], by $18.8\%$ and the premier ViT-based model, PSKD [55], by $19.2\%$ in the mAP@all metric. In comparison with CLIP-based models, SpLIP achieves an improvement over MARL by $1.5\%$ in mAP values, with analogous enhancements observed across all datasets, indicating at least a $2\%$ boost in mAP scores. In Fig. 3, we show qualitative comparisons between our proposed SpLIP and [39] on ZS-SBIR, highlighting our improved retrieval results for different categories.
Further insights into the performance on more stringent GZS-SBIR tasks are provided in Table 4.1, where SpLIP demonstrates its ability to alleviate model bias towards training classes significantly. It outmatches the second-best approach by $4.8\%$ on Sketchy and $4.1\%$ on TU-Berlin in mAP scores, unequivocally evidencing enhanced generalization capabilities.
@@
Methods Sketchy-Ext TU-Berlin-Ext
@@
Table 4: Comparison of ZS-SBIR across datasets while training with the Sketchy extended dataset and tested on TU-Berlin extended and Quickdraw extended datasets. represents the results reproduced by us.
@@
- Across dataset ZS-SBIR: Leveraging on the across-dataset generalizability capabilities of CLIP, we assess the effectiveness of our proposed SpLIP on the ZS-SBIR across dataset setting by [26], where the model is trained on the Sketchy extended dataset and evaluated on 21 unseen classes of TU-Berlin extended and 11 unseen classes of Quickdraw extended datasets, respectively. In Table 10, our findings demonstrate that SpLIP outperforms CLIP-AT [39] in terms of mAP@all and P@100 metrics by $14.2\%$ and $12.9\%$ on the TU-Berlin extended dataset, and by $15.1\%$ and $13.6\%$ on the Quickdraw extended dataset.
@@ 4.2 Ablation analysis
Besides the following, ablations on the number of training samples, the ratio of known and unknown classes etc. are mentioned in the Supplementary.
- Analysis of the loss components: In Table 6, we explore the influence of various loss components as described in Section 3.5 on the ZS-SBIR and FG-ZS-SBIR tasks, utilizing the Sketchy dataset for evaluation. Incorporation of $\mathcal{L}_{triplet}$ demonstrates a notable improvement in performance, which is further sharply enhanced by $\mathcal{L}_{class}$. The addition of $\mathcal{L}_{cjs}$ contributes to an approximate $3-4\%$ increase in results for both tasks. Specifically, we find that both the components of $\mathcal{L}_{cjs}$ show improvements, i.e., using $\mathcal{L}_{class}+\mathcal{L}_{triplet}+\mathcal{L}_{margin}$ improves the results of $\mathcal{L}_{class}+\mathcal{L}_{triplet}$ by $\approx 1-2\%$, and the use of full $\mathcal{L}_{cjs}$ offers further improvements by $2\%$ for ZS-SBIR and $\approx 2-3\%$ for FG-ZS-SBIR.
Figure 3: Qualitative comparisons between SpLIP and [39]. Improved retrieval outcomes for more ambiguous classes can be observed with SpLIP.
- Ablation with learnable blocks of SpLIP: In Table 6, we elucidate the significance of each learnable module within SpLIP, $\mathcal{B}_{v}$, $\mathcal{B}_{t}$, $\mathcal{B}_{vt}$, and LayerNorm parameters, for both ZS-SBIR and FG-ZS-SBIR tasks. Note that the exclusion of one of these blocks means the respective prompt token sharing is disabled. Consistent with findings from [39], rendering the parameters of LayerNorm modules learnable results in a performance boost of approximately $2\%$. Conversely, the absence of cross-modal token sharing, whether in one or both directions, detrimentally impacts outcomes. Specifically, we note a decline in precision by at least $4\%$ for ZS-SBIR and $3\%$ for FG-ZS-SBIR tasks, underscoring the critical role of token sharing in enhancing model efficacy across both retrieval challenges.
- Shallow to deep prompting in SpLIP: Additionally, we assess the effect of employing deep prompting across separate textual and visual branches, alongside our integrated multi-modal approach, as depicted in Figure 4 (Left). This evaluation reveals a general trend where mAP values ascend with the inclusion of more layers from $\mathcal{F}_{v}$/$\mathcal{F}_{t}$ for all three scenarios. Notably, the enhancements achieved through multi-modal prompting distinctly surpass those of the uni-modal counterparts, emphasizing the superior efficacy of leveraging both textual and visual cues in a cohesive manner.
@@
Table 5: Ablation of loss terms (Equation 11) on the Sketchy-ext dataset for both the ZS-SBIR and FG-ZS-SBIR tasks.
@@
Figure 4: Left: Effects of including more layers into our proposed deep prompting. We show the effects of prompting individually for the vision and textual branches of CLIP, followed by the proposed multi-modal prompting. Right: B1 refers effects of replacing the multi-modal prompting of ours by that of [20], B2 refers to replacing the conditional cross-modal jigsaw solver by the patch permutation-based triplet objective proposed in [39], and B3 denotes SpLIP with a fixed $\mu$ of $0.2$ for $\mathcal{L}_{triplet}$.
@@ 5 Takeaways
This paper unveils SpLIP, a novel CLIP-based framework tailored for ZS, GZS, and FG-ZS SBIR tasks. Our methodology is distinguished by three innovations: a multi-modal prompt learning strategy enhancing cross-modal knowledge sharing and embedding learning within text-sketch-photo triads; the use of CLIP’s textual class-name embeddings to dynamically adjust margins for the sketch-photo triplet loss; and a unique conditional cross-modal jigsaw task designed to fine-tune patch-level sketch-photo associations, by repurposing the stand-alone jigsaw task with a metric objective. SpLIP’s efficacy is validated across various benchmarks, showcasing its dominance in all tasks consistently. Our future work is deemed to extend to broader vision-language domains, enriching our understanding of visual semantics.
@@ Appendix 0.B Dataset descriptions
In this work, we evaluate the SpLIP method’s effectiveness across leading ZS-SBIR datasets, including Sketchy Extended, TU-Berlin Extended, and QuickDraw Extended.
Sketchy Extended [27]: This dataset enlarges the original Sketchy collection [44] to 73,002 images across 125 categories, with an average of 604 sketches and 584 images per category. For zero-shot analysis, two distinct splits are employed. In split-1 [11], 25 categories are randomly chosen for testing, leaving the remaining 100 for training. Split-2 [57] isolates 21 classes not present in the ImageNet [38] dataset for testing, with the other 104 classes designated for training.
TU-Berlin Extended [61]: An augmentation of the original TU-Berlin dataset [13], which included 20,000 sketches over 250 categories. This extension incorporates 204,489 natural images [60], averaging 787 images per category, though with notable class-wise image imbalances. Adhering to the partitioning protocol from [46], we allocate 30 classes for testing and 220 for training, ensuring each test class has a minimum of 400 images to mitigate the imbalance.
QuickDraw Extended [7]: A large-scale dataset designed for ZS-SBIR, extending the Google QuickDraw dataset [17] with detailed photographs. It includes 110 categories, featuring 330,000 sketches (3,000 per category) and 204,000 Flickr-sourced images, each tagged appropriately. The partitioning strategy from [7] is used, segregating 30 test classes not found in ImageNet, with the remaining 80 classes utilized for training.
@@ Appendix 0.C List of important variables and their significance
The findings from these ablation studies highlight a trend where an increase in the number of tokens correlates with improved performance. Specifically, the configuration with tokens set to (4, 4, 2) achieved superior results for FG-ZS-SBIR tasks on the Sketchy-Extended dataset. Moreover, it was observed that employing a higher number of prompts tends to diminish the raw feature representation, affecting the overall performance of the system. This observation underscores the balance required between the number of prompts and the preservation of feature quality for optimal performance.

### v2

@@ 3.5 Loss functions and inference
$\mathcal{L}_{triplet}$: Cross-visual modality triplet loss with proposed adaptive margin: Given the triplet of embeddings $(\mathcal{F}_{v}(s_{a}),\mathcal{F}_{v}(p_{a}^{+}),\mathcal{F}_{v}(p_{a}^{-}))$, we formulate a triplet objective aimed at minimizing the distance between $\mathcal{F}_{v}(s_{a})$ and $\mathcal{F}_{v}(p_{a}^{+})$ while maximizing the distance between $\mathcal{F}_{v}(s_{a})$ and $\mathcal{F}_{v}(p_{a}^{-})$. Traditional triplet objectives employ a fixed margin, which may not be optimal for zero-shot tasks. In contrast, we propose leveraging the semantic space of CLIP to define the margin, utilizing the embeddings of positive and negative class names from $\mathcal{F}_{t}$. Let $\mathcal{F}_{t}(c^{+})$ and $\mathcal{F}_{t}(c^{-})$ represent the semantic class embeddings for $p_{a}^{+}/s_{a}$ and $p_{a}^{-}$, respectively. We define $\mu(c^{+},c^{-})=\cos(\mathcal{F}_{t}(c^{+}),\mathcal{F}_{t}(c^{-}))$. Consequently, the margin increases when $c^{+}$ and $c^{-}$ are semantically close. The loss is,
@@
$\mathcal{L}_{class}$: Text-image classification loss: A supervised contrastive loss is introduced to correctly classify sketches and photos in $\mathcal{G}^{s}$ based on the class-wise text prompts outlined in Section 3.1. In Eq. 7, $(\mathbf{I,\mathbf{Y}})$ denotes a sketch or a photo along with the associated one-hot label vector: $\mathbf{I}\in\{\mathcal{P}^{s},\mathcal{S}^{s}\}$, and $\mathbf{Y}=[y^{1},\cdots,y^{|\mathcal{C}^{s}|}]$.
@@ 4 Experiments and Results
- Datasets: We evaluate SpLIP on three benchmark datasets for categorical (G)ZS-SBIR: Sketchy-Ext [11, 57], TU-Berlin-Ext [27], and QuickDraw-Ext [7], following the established training-validation protocols [5, 11]. For FG-ZS-SBIR, which necessitates precise sketch-photo matching, we incorporate the Sketchy dataset [44]. Further details, including the splits of Sketchy-1-Ext [11] & Sketchy-2-Ext [57] are described in the Supplementary.
- Implementation details, training and evaluation protocols: For $\mathcal{F}_{v}$, we select the ViT-B/32 backbone of CLIP, while the Transformer-based text encoder is considered for $\mathcal{F}_{t}$. Besides, $\mathcal{B}_{t}$ employs three linear layers to convert $\mathbf{E}_{0}$ into $m=4$ learnable textual tokens. In contrast, $\mathcal{B}_{v}$ utilizes a singular linear layer to adapt to the visual dimension, producing four learnable visual tokens for a matching batch size. Meanwhile, $\mathcal{B}_{vt}$ incorporates linear layers and employs a bottleneck architecture consisting of two layers (Linear-ReLU-Linear) for the creation of $n=2$ layer-specific visual tokens. Additionally, $\mathcal{F}_{js}$ is designed with two encoder layers, followed by a classifier, to accurately decode the patch arrangements of $s^{\prime}_{a}$ given ($p_{a}^{+}$/$p_{a}^{-}$/$s_{a}$).
The training process spans 60 epochs, initiating with a warm-up learning rate of $0.001$ and utilizing the Adam optimizer [22] alongside a scheduler. The batch size is configured to $192$ for both Sketchy-Ext and TU-Berlin-Ext datasets, whereas a batch size of $64$ is adopted for QuickDraw-Ext. $\alpha$ and $\beta$ are fixed through grid search.
@@
Following the literature [7, 55], our evaluation for ZS-SBIR considers the top $200$ retrieved photos, where we report the mean Average Precision score (mAP@all) and precision at $200$ (P@200). Aligning with the recent trends, however, we specifically report precision at $100$ (P@100) for the TU-Berlin-Ext dataset and mAP at $200$ (mAP@200) for the Sketchy dataset. For FG-ZS-SBIR, accuracy is evaluated by considering only a single category at a time [33], denoted as Acc@$\mathcal{K}$ for Sketchy. This metric reflects the percentage of sketches that have their true matched photo within the top-$\mathcal{K}$ list, with our focus being on the Top-1 and Top-5 accuracy metrics [39].
@@ 4.1 Comparison to the literature
- (G)ZS-SBIR and FG-ZS-SBIR: Table 1 provides a comparative analysis of SpLIP against methods utilizing ConvNet, ViT, and CLIP backbones across the extended versions of the two splits of the Sketchy, TU-Berlin, and QuickDraw datasets, respectively in categorical ZS-SBIR. Remarkably, SpLIP outperforms all models based on ConvNet and ViT backbones by a substantial margin across all metrics. For instance, on the challenging QuickDraw-Ext dataset, SpLIP surpasses the leading ConvNet-based method, BDA [5], by $18.8\%$ and the premier ViT-based model, PSKD [55], by $19.2\%$ in the mAP@all metric. In comparison with CLIP-based models, SpLIP achieves an improvement over MARL by $1.5\%$ in mAP values, with analogous enhancements observed across all datasets, indicating at least a $2\%$ boost in mAP scores. In Fig. 3, we show qualitative comparisons between our proposed SpLIP and [39] on ZS-SBIR, highlighting our improved retrieval results for different categories.
Further insights into the performance on more stringent GZS-SBIR tasks are provided in Table 4.1, where SpLIP significantly demonstrates its ability to alleviate model bias towards training classes. It outmatches the second-best approach by $4.8\%$ on Sketchy-Ext and $4.1\%$ on TU-Berlin-Ext in mAP scores, unequivocally evidencing enhanced generalization capabilities.
@@
Methods Sketchy-2-Ext[57] TU-Berlin-Ext[27]
@@
Table 4: Comparison of ZS-SBIR across datasets while training with the Sketchy-Ext [57] dataset and tested on TU-Berlin-Ext and Quickdraw-Ext datasets. represents the results reproduced by us.
@@
- Across dataset ZS-SBIR: Leveraging on the across-dataset generalizability capabilities of CLIP, we assess the effectiveness of our proposed SpLIP on the ZS-SBIR across dataset setting by [26], where the model is trained on the Sketchy-Ext dataset and evaluated on 21 unseen classes of TU-Berlin-Ext and 11 unseen classes of Quickdraw-Ext datasets, respectively. In Table 10, our findings demonstrate that SpLIP outperforms CLIP-AT [39] in terms of mAP@all and P@100 metrics by $14.2\%$ and $12.9\%$ on the TU-Berlin-Ext dataset, and by $15.1\%$ and $13.6\%$ on the Quickdraw-Ext dataset.
@@ 4.2 Ablation analysis
Besides the following, ablations on the number of training samples, the ratio of known and unknown classes, etc., are mentioned in the Supplementary.
- Analysis of the loss components: In Table 6, we explore the influence of various loss components as described in Section 3.5 on the ZS-SBIR and FG-ZS-SBIR tasks, utilizing the Sketchy dataset for evaluation. Incorporation of $\mathcal{L}_{triplet}$ demonstrates a notable performance improvement, which is further sharply enhanced by $\mathcal{L}_{class}$. Adding $\mathcal{L}_{cjs}$ contributes to an approximate $3-4\%$ increase in results for both tasks. Specifically, we find that both the components of $\mathcal{L}_{cjs}$ show improvements, i.e., using $\mathcal{L}_{class}+\mathcal{L}_{triplet}+\mathcal{L}_{margin}$ improves the results of $\mathcal{L}_{class}+\mathcal{L}_{triplet}$ by $\approx 1-2\%$, and the use of full $\mathcal{L}_{cjs}$ offers further improvements by $2\%$ for ZS-SBIR and $\approx 2-3\%$ for FG-ZS-SBIR.
Figure 3: Qualitative comparisons between SpLIP and [39] in categorical ZS-SBIR. Improved retrieval outcomes for more ambiguous classes can be observed with SpLIP.
- Ablation with learnable blocks of SpLIP: In Table 6, we elucidate the significance of each learnable module within SpLIP, $\mathcal{B}_{v}$, $\mathcal{B}_{t}$, $\mathcal{B}_{vt}$, and LayerNorm parameters, for both ZS-SBIR and FG-ZS-SBIR tasks. Note that excluding one of these blocks means the respective prompt token sharing is disabled. Consistent with findings from [39], rendering the parameters of LayerNorm modules learnable results in a performance boost of approximately $2\%$. Conversely, the absence of cross-modal token sharing detrimentally impacts outcomes, whether in one or both directions. Specifically, we note a decline in precision by at least $4\%$ for ZS-SBIR and $3\%$ for FG-ZS-SBIR tasks, underscoring the critical role of token sharing in enhancing model efficacy across both retrieval challenges.
- Shallow to deep prompting in SpLIP: Additionally, we assess the effect of employing deep prompting across separate textual and visual branches, alongside our integrated multi-modal approach, as depicted in Figure 4 (Left). This evaluation reveals a general trend where mAP values ascend with the inclusion of more layers from $\mathcal{F}_{v}$/$\mathcal{F}_{t}$ for all three scenarios. Notably, the enhancements achieved through multi-modal prompting distinctly surpass those of the uni-modal counterparts, emphasizing the superior efficacy of cohesively leveraging both textual and visual cues.
@@
Table 5: Ablation of loss terms (Equation 11) on the Sketchy dataset.
@@
Figure 4: Left: Effects of including more layers into our proposed deep prompting. We show the effects of prompting individually for CLIP’s vision and textual branches, followed by the proposed multi-modal prompting. Right: B1 refers effects of replacing the multi-modal prompting of ours by that of [20], B2 refers to replacing the conditional cross-modal jigsaw solver by the patch permutation-based triplet objective proposed in [39], and B3 denotes SpLIP with a fixed $\mu$ of $0.2$ for $\mathcal{L}_{triplet}$.
@@ 5 Takeaways
This paper unveils SpLIP, a novel CLIP-based framework tailored for ZS, GZS, and FG-ZS SBIR tasks. Our methodology is distinguished by three innovations: a multi-modal prompt learning strategy enhancing cross-modal knowledge sharing and embedding learning within text-sketch-photo triads; the use of CLIP’s textual class-name embeddings to dynamically adjust margins for the sketch-photo triplet loss; and a unique conditional cross-modal jigsaw task designed to fine-tune patch-level sketch-photo associations, by repurposing the stand-alone jigsaw task with a metric objective. SpLIP’s efficacy is validated across various benchmarks, consistently showcasing its dominance in all tasks. Our future work will extend to broader vision-language domains, enriching our understanding of visual semantics. The authors sincerely acknowledge the tremendous support from AWL Inc, Japan.
@@ Appendix 0.B Dataset descriptions
In this work, we evaluate the SpLIP method’s effectiveness across leading ZS-SBIR datasets, including Sketchy-Ext, TU-Berlin-Ext, and QuickDraw-Ext.
Sketchy-Ext [11, 57]: This dataset enlarges the original Sketchy collection [44] to 73,002 images across 125 categories, with an average of 604 sketches and 584 images per category. For zero-shot analysis, two distinct splits are employed. In split-1, i.e., Sketchy-1-Ext [11], 25 categories are randomly chosen for testing, leaving the remaining 100 for training. On the other hand, Sketchy-2-Ext [57] isolates 21 classes not present in the ImageNet [38] dataset for testing, with the other 104 classes designated for training.
TU-Berlin-Ext [27]: An augmentation of the original TU-Berlin [13] dataset, which included 20,000 sketches over 250 categories. This extension incorporates 204,489 natural images [60], averaging 787 images per category, though with notable class-wise image imbalances. Adhering to the partitioning protocol from [46], we allocate 30 classes for testing and 220 for training, ensuring each test class has a minimum of 400 images to mitigate the imbalance.
QuickDraw-Ext [7]: A large-scale dataset designed for ZS-SBIR, extending the Google QuickDraw dataset [17] with detailed photographs. It includes 110 categories, featuring 330,000 sketches (3,000 per category) and 204,000 Flickr-sourced images, each tagged appropriately. The partitioning strategy from [7] is used, segregating 30 test classes not found in ImageNet, with the remaining 80 classes utilized for training.
@@ Appendix 0.C List of important variables and their significance
The findings from these ablation studies highlight a trend where an increase in the number of tokens correlates with improved performance. Specifically, the configuration with tokens set to (4, 4, 2) achieved superior results for FG-ZS-SBIR tasks on the Sketchy-Ext dataset. Moreover, it was observed that employing a higher number of prompts tends to diminish the raw feature representation, affecting the overall performance of the system. This observation underscores the balance required between the number of prompts and the preservation of feature quality for optimal performance.